OpenStudy (anonymous):

This OnE i Don't know Anything about it =..= i was Absent that DaY .. So HElp me With it (--)

OpenStudy (anonymous):

here is the Question .. let \[f(x)= x ^{3}+4x +3 \] Evaluate \[\frac{ d }{ dx} (f ^{-1} (3))\]

OpenStudy (anonymous):

are you sure that this is the question ? f-1(3) is a number hence constant so the derivative of this is 0.

OpenStudy (anonymous):

f' = 3x^2 + 4 f'(3) = 3*9 +4 = 31

OpenStudy (anonymous):

or you need d/dx ( f-1(x)) at the point 3 ?

OpenStudy (anonymous):

you do not need to invert the function to find the derivative of the inverse at a given number

OpenStudy (anonymous):

one thing that should be clear from the start is that \(f(0)=3\) right? it is the constant therefore you know \(f^{-1}(3)=0\) we need that number

OpenStudy (anonymous):

(×_× )

OpenStudy (anonymous):

Some oNe tell me what is the first SteP ? <>_<>

OpenStudy (anonymous):

i wrote the first step. the first step is figuring out what \(f^{-1}(3)\) is

OpenStudy (anonymous):

how -.-

OpenStudy (anonymous):

you pretty much have to make a good guess for this one, but it was easy to find

OpenStudy (anonymous):

i thought "i must find \(f^{-1}(3)\)" meaning "i must think what i would plug in to get 3 out" and there was the constant 3 staring me in the face, so i reasoned that \(f(0)=3\) for sure

OpenStudy (anonymous):

so now we put aside the fact that we know \(f^{-1}(3)=0\) and then see how to find the derivative of an inverse function

OpenStudy (anonymous):

(-_-) what r u talking about

OpenStudy (anonymous):

i was describing in the best way i know how, the method i used to find \(f^{-1}(3)\) you need that number, as you can see from the question

OpenStudy (anonymous):

without it, we cannot start

OpenStudy (anonymous):

OMg .. i Feel like im the dumpiest Person on earth (º_º')

OpenStudy (anonymous):

i can write an equation if you like we can solve \[x^3+4x+3=3\] for \(x\) and see what we get

OpenStudy (anonymous):

step one \[x^3+4x=0\] step two \[x(x^2+4)=0\] and therefore \(x=0\)

OpenStudy (anonymous):

should i take the derivative for the function f(x) first ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

(π_π)

OpenStudy (anonymous):

lets make it clear what the steps are 1) first find \(f^{-1}(3)\) 2) then find an expression for \(\frac{d}{dx}[f^{-1}(x)]\) in terms of \(f\), \(f'\) and \(f^{-1}\) 3) then evaluate i will write the steps and maybe it will be clear

OpenStudy (anonymous):

step one is done, we know \(f^{-1}(3)=0\) so far so good?

OpenStudy (anonymous):

from where we got Zer0

OpenStudy (anonymous):

u mean u found the value for x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but there will be more than one value for x

OpenStudy (anonymous):

i found the value of \(x\) that would make \(f(x)=3\)

OpenStudy (anonymous):

ohh yeah ur right

OpenStudy (anonymous):

if there is, then the function is not invertible (has no inverse) but no, there is only one value

OpenStudy (anonymous):

yeah yeah ur right :B

OpenStudy (anonymous):

ok so far so good then? we know \(f(0)=3\) and so \(f^{-1}(3)=0\) yes?

OpenStudy (anonymous):

ummm okie So far so Good :D

OpenStudy (anonymous):

ok good. now the following is always true \[\frac{d}{dx}[f^{-1}(x)]=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

noo ur wrong >:[ wait its not equal to zero =..=

OpenStudy (anonymous):

i am positive that \(f^{-1}(3)=0\), and i know that because \(f(0)=3\)

OpenStudy (anonymous):

what makes you think it is not zero?

OpenStudy (anonymous):

>_ >'

OpenStudy (anonymous):

okie whatever u say professor =.=

OpenStudy (anonymous):

its Zero cuz it is ZerO

OpenStudy (anonymous):

▼_▼

OpenStudy (anonymous):

so we need \[\frac{1}{f'(f^{-1}(3))}\] for for the answer, in other words \[\frac{1}{f'(0)}\]

OpenStudy (anonymous):

that is why it was important that we find \(f^{-1}(3)\) first. without it we cannot compute this number with it, it is easy enough, since the derivative of \(f(x)=x^3+4x+3\) is \(f'(x)=3x^2+4\) and so \(f'(0)=4\) making \(\frac{d}{dx}[f^{-1}(3)]=\frac{1}{4}\)

OpenStudy (anonymous):

to do this problem you need to know two facts \[f^{-1}(3)=0\] and \[\frac{d}{dx}[f^{-1}(x)]=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

>_ > im Studying ur comments Sir .. i need TiMe ●,●゚

OpenStudy (anonymous):

take the time, all the steps are there. the only thing i did not write is why it is true that \[\frac{d}{dx}[f^{-1}(x)]=\frac{1}{f'(f^{-1}(x))}\] i can explain it if you like, it is a direct consequence of the chain rule, but in order to do this problem you need to use this fact (not necessarily understand where it comes from)

OpenStudy (anonymous):

So the final Answer iS just 1/4 ? ._.

OpenStudy (anonymous):

yes, we found \(f^{-1}(3)=0\) we found \(f'(x)=3x^2+4\) we found \(f'(0)=4\) and we know the reciprocal of \(4\) is \(\frac{1}{4}\)

OpenStudy (anonymous):

yeah i will see how u got the answer ^_^' (..;) anyways .... Thanks Sir @satellite73 sankiyo very much X3 OAO

OpenStudy (anonymous):

yw